### v4

@@ Abstract
Factorization machine (FM) is a prevalent approach to modeling pairwise (second-order) feature interactions when dealing with high-dimensional sparse data. However, on the one hand, FM fails to capture higher-order feature interactions suffering from combinatorial expansion. On the other hand, taking into account interactions between every pair of features may introduce noise and degrade prediction accuracy. To solve the problems, we propose a novel approach, Graph Factorization Machine (GraphFM), by naturally representing features in the graph structure. In particular, we design a mechanism to select the beneficial feature interactions and formulate them as edges between features. Then the proposed model, which integrates the interaction function of FM into the feature aggregation strategy of Graph Neural Network (GNN), can model arbitrary-order feature interactions on the graph-structured features by stacking layers. Experimental results on several real-world datasets have demonstrated the rationality and effectiveness of our proposed approach. The code and data are available at https://github.com/CRIPAC-DIG/GraphCTR.
@@ 1 Introduction
Predictive analytics is a fundamental task in machine learning (ML) and data mining (DM), which involves using input features to predict an output target, such as a real value for regression or categorical labels for classification. This is particularly important for web applications, such as online advertising and recommender systems He et al. (2017); He and Chua (2017). Distinct from continuous features which can be naturally found in images and audio, the features for web applications are mostly sparse and categorical. To accurately perform predictive analytics on these types of features, it is important to consider the interactions between them. As an example, consider a scenario in which we want to predict users’ preferences for movies based on five categorical variables: (1) Language = {English, Chinese, Japanese, … }, (2) Genre = {action, fiction, … }, (3) Director = {Ang Lee, Christopher Nolan, … }, (4) Stars = {Bruce Lee, Leonardo DiCaprio, … } and (5) Release Date = {1991, 1992, … }. To capture the impact of these feature interactions, a model might consider a 3-order cross feature such as (Genre = fiction, Director = Christopher Nolan, Starring = Leonardo DiCaprio) or (Language = Chinese, Genre = action, Starring = Bruce Lee) as potentially indicating higher user preferences.
@@
Factorization machine (FM) Rendle (2010, 2012) are a popular and effective method for modeling feature interactions, which involve learning a latent vector for each one-hot encoded feature and modeling the pairwise (second-order) interactions between them through the inner product of their respective vectors. FM has been widely used in the field of recommender systems and click-through rate predictions due to its simplicity and effectiveness. However, because FM considers all feature interactions, it has two main drawbacks.
@@
One of the main limitations of FM is that it is not able to capture higher-order feature interactions, which are interactions between three or more features. While higher-order FM (HOFM) has been proposed Rendle (2010, 2012) as a way to address this issue, it suffers from high complexity due to the combinatorial expansion of higher-order interactions. This makes HOFM difficult to use in practice. To address the limitations of FM in capturing higher-order feature interactions, several variants have been proposed that utilize deep neural networks (DNNs) Zhang et al. (2016); He and Chua (2017); Cheng et al. (2016); Guo et al. (2017). Factorisation-machine supported Neural Networks (FNNs) Zhang et al. (2016) apply DNN on top of pre-trained factorization machines to model high-order interactions. Neural Factorization Machines (NFM) He and Chua (2017) design a bi-interaction layer to learn the pairwise feature interaction and apply DNN to learn the higher-order ones. Wide&Deep Cheng et al. (2016) introduces a hybrid architecture containing both shallow and deep components to jointly learn low-order and high-order feature interactions. DeepFM Guo et al. (2017) similarly combines a shallow component with a deep one to learn both types of interactions. While these DNN-based models can effectively learn high-order feature interactions, they do so in an implicit, bit-wise manner. Consequently, they may lack the ability to provide persuasive rationales for their outputs.
@@
In addition to not being able to effectively capture higher-order feature interactions, FM is also suboptimal because it considers the interactions between every pair of features, even if some of these interactions may not be beneficial for prediction Zhang et al. (2016); Su et al. (2020). These unhelpful feature interactions can introduce noise and lead to overfitting, as they do not provide useful information but make it harder to train the model. For example, in the context of predicting movie preferences, the feature interactions between Language and Release Date might not be relevant and, therefore, not provide useful information for prediction. Ignoring these irrelevant feature interactions can improve model training. To solve this problem, the Attentional Decomposition Machine (AFM) Xiao et al. (2017) distinguishes the importance of the factorized interaction by reweighing each cross-feature using the attentional score Bahdanau et al. (2014), i.e. the influence of useless cross-features is reduced by assigning lower weights. However, it requires a predefined maximum order, which limits the potential of the model to find discriminative crossing features. Therefore, the Adaptive Factorization Network (AFN) Cheng et al. (2020) uses a logarithmic neural transformation layer composed of multiple vector-wise logarithmic neurons to automatically learn the powers (i.e. the order) of features in a potentially useful combination, thereby adaptively learning cross-features and their weights from the data.
@@
Currently, Graph Neural Networks (GNN) Kipf and Welling (2017); Hamilton et al. (2017); Veličković et al. (2018) have recently emerged as an effective class of models for capturing high-order relationships between nodes in a graph and have achieved state-of-the-art results on a variety of tasks such as computer vision Li et al. (2017), neural language processing Marcheggiani and Titov (2017); Yao et al. (2019), and recommender systems Wang et al. (2019); Wu et al. (2019). At their core, GNNs learn node embeddings by iteratively aggregating features from the neighboring nodes, layer by layer. This allows them to explicitly encode high-order relationships between nodes in the embeddings. GNNs have shown great potential for modeling high-order feature interactions for click-through rate prediction. Fi-GNN Li et al. (2019) proposes to connect each pair of features and treat the multi-field features as a fully-connected graph, using a Gated Graph Neural Network (GGNN) Li et al. (2015) to model feature interactions on the graph. Graph Factorizer Machine (GFM) Xi et al. (2020) utilizes FM to aggregate second-order neighbour messages, and utilizes the superposition of multiple GFM layers to aggregate higher-order neighbour messages to achieve multi-order interaction from neighborhood for recommendation. Graph-Convolved Factorization Machines (GCFM) Zheng et al. (2021) developed the Graph-Convolved Feature Crossing (GCFC) layer to traverse all features for each input example and leveraged the features of each sample to compute the corresponding multi-feature interaction graph and propagated its influence on other features. KD-DAGFM Tian et al. (2023) proposes a directed acyclic graph based model, which can be aligned with the DP Dudzik and Veličković (2022) algorithm to improve knowledge distillation (KD)Hinton et al. (2015) capability. However, not all feature interactions are beneficial, and GNNs rely on the assumption that neighboring nodes share similar features, which may not always hold in the context of feature interaction modeling.
@@
In summary, when dealing with feature interactions, FM suffers intrinsic drawbacks. We thus propose a novel model Graph Factorization Machine (GraphFM), which takes advantage of GNN to overcome the problems of FM for feature interaction modeling. By treating features as nodes and feature interactions as the edges between them, the selected beneficial feature interactions can be viewed as a graph. We thus devise a novel technique to select the beneficial feature interactions, which is also to infer the graph structure. Then we adopt an attentional aggregation strategy to aggregate these selected beneficial interactions to update the feature representations. Specifically, to accommodate the polysemy of feature interactions in different semantic spaces, we utilize a multi-head attention mechanism Vaswani et al. (2017); Veličković et al. (2018). Each layer of our proposed model produces higher-order interactions based on the existing ones and thus the highest-order of interactions is determined by layer depth. Since our proposed approach selects the beneficial feature interactions and models them in an explicit manner, it has high efficiency in analyzing high-order feature interactions and thus provides rationales for the model outcome. Through extensive experiments conducted on CTR benchmark and recommender system datasets, we verify the rationality, effectiveness, and interpretability of our proposed approach.
@@
Overall, the main contributions of this work are threefold: (1) We analyze the shortcomings and strengths of FM and GNN in modeling feature interactions. To solve their problems and leverage strengths, we propose a novel model GraphFM for feature interaction modeling. (2) By treating features as nodes and their pairwise feature interactions as edges, we bridge the gap between GNN and FM, and make it feasible to leverage the strength of GNN to solve the problem of FM. (3) Extensive experiments are conducted on CTR benchmark and recommender system datasets to evaluate the effectiveness and interpretability of our proposed method. We show that GraphFM can provide persuasive rationales for the feature interaction modeling and prediction-making process.
@@ 2 Related Work
In this work, we proposed a graph neural network-based approach to modeling feature interactions. We design a feature interaction selection mechanism, which can be seen as learning the graph structure by viewing the feature interactions as edges between features. In this section, we review three lines of research that are relevant to this work: 1) techniques for learning feature interactions, 2) Graph Neural Networks, and 3) graph structure learning methods.
2.1 Feature Interaction Modeling
@@
Modeling feature interactions is a crucial aspect of predictive analytics and has been widely studied in the literature. FM Rendle (2010) is a popular method that learns pairwise feature interactions through vector inner products. Since its introduction, several variants of FM have been proposed, including Field-aware factorization machine (FFM) Juan et al. (2016) which takes into account field information and introduces field-aware embeddings, and AFM Xiao et al. (2017), which considers the weight of different second-order feature interactions. FmFM Sun et al. (2021) model the interactions of field pairs as a matrix and utilizes kernel product to capture field interaction. However, these approaches are limited to modeling second-order interactions, which may not be sufficient in some cases.
@@
As deep neural networks (DNNs) have proven successful in a variety of fields, researchers have begun using them to learn high-order feature interactions due to their deeper structures and nonlinear activation functions. The general approach is to concatenate the representations of different feature fields and feed them into a DNN to learn the high-order feature interactions. Factorization-machine supported Neural Networks (FNNs) Zhang et al. (2016) use pre-trained factorization machines to create field embeddings before applying a DNN, while Product-based Neural Networks (PNNs) Qu et al. (2016) model both second-order and high-order interactions through the use of a product layer between the field embedding layer and the DNN layer. Like PNNs, Neural Factorization Machines (NFMs) He and Chua (2017) also use a separate layer to model second-order interactions, but they use a Bi-Interaction Pooling layer instead of a product layer and follow it with summation rather than concatenation. Other approaches to modeling second-order and high-order interactions jointly use hybrid architectures. The Wide&Deep Cheng et al. (2016) and DeepFM Guo et al. (2017) contain a wide part modeling the low-order interaction and a deep part modeling the high-order interaction. However, like the other DNN-based approaches, these models learn high-order feature interactions in an implicit, bit-wise manner and may lack transparency in their feature interaction modeling process and model outputs. As a result, some studies have attempted to learn feature interactions in an explicit fashion through the use of specifically designed networks. Deep&Cross Wang et al. (2017) introduces a CrossNet that takes the outer product of features at the bit level, while xDeepFM Lian et al. (2018) uses a Compressed Interaction Network(CIN) to take the outer product at the vector level and then compresses the resulting feature maps to update the feature representations. However, xDeepFM has been found to have issues with generalization and scalability, and it has relatively high complexity due to its consideration of all pairwise bit-level interactions. DCNV2 Wang et al. (2021) similarly uses CIN to learn efficient explicit and implicit feature intersections, but it additionally leverages low-rank techniques to approximate feature crosses in subspace for better performance and latency trade-offs.
@@
More recently, some studies have attempted to use attention mechanisms to model feature interactions in a more interpretable way. HoAFM Tao et al. (2019) updates feature representations by attentively aggregating the representations of co-occurring features, while AutoInt Song et al. (2019) uses a multi-head self-attention mechanism to explicitly model feature interactions. InterHAt Li et al. (2020) is another model that uses an attentional aggregation strategy with residual connections to learn feature representations and model feature interactions. However, even with the use of attention mechanisms to account for the weight of each pair of feature interactions, aggregating all interactions together can still introduce noise and degrade prediction accuracy. To address these issues, some recent studies have attempted to identify beneficial feature interactions automatically. AutoFIS Liu et al. (2020) is a two-stage algorithm that uses a gate operation to search and model beneficial feature interactions, but there is a loss of information between the stages, and the modeling process is not interpretable. AFN Cheng et al. (2020) uses a logarithmic neural network to adaptively learn high-order feature interactions, and SIGN Su et al. (2020) utilizes mutual information to detect beneficial feature interactions and a linear aggregation strategy to model them. However, these approaches may not be expressive or interpretable enough.
@@ 2.2 Graph Neural Networks
The graph is a kind of data structure that reflects a set of entities (nodes) and their relations (edges). Graph neural networks (GNNs), as deep learning architectures on graph-structured data, have attracted increasing attention. The concept of GNNs is first proposed by Gori et al. (2005), and further elaborated in Scarselli et al. (2009). Nowadays, most of the prevailing GNN models follow the neighborhood aggregation strategy, that is, to learn the latent node representations via aggregating the features of neighborhoods layer by layer. The high-order relations between nodes can be modeled explicitly by stacking layers. Gated Graph Neural Networks (GGNN) Li et al. (2015) uses GRU Cho et al. (2014) to update the node representations based on the aggregated neighborhood feature information. Though based on graph spectral theory Bruna et al. (2013), the learning process of graph convolutional networks (GCN) Kipf and Welling (2017) also can be considered as a mean-pooling neighborhood aggregation. GraphSAGE Hamilton et al. (2017) concatenates the node features and introduces three mean/max/LSTM aggregators to pool the neighborhood information. Graph attention network (GAT) Veličković et al. (2018) incorporates the attention mechanism to measure the weights of neighbors when aggregating neighborhood information of a node.
@@
Due to the strength in modeling relations on graph-structured data, GNN has been widely applied to various applications like neural machine translation Beck et al. (2018), semantic segmentation Qi et al. (2017), image classification Marino et al. (2017), situation recognition Li et al. (2017), recommendation Wu et al. (2019); Chen et al. (2020); Wang et al. (2019), script event prediction Li et al. (2018), and fashion analysis Cui et al. (2019). Fi-GNN Li et al. (2019) is the first attempt to exploit GNN for feature interaction modeling. It first proposes to connect all the feature fields, and thus the multi-field features can be treated as a fully-connected graph. Then it utilizes GGNN Li et al. (2015) to model high-order feature interactions on the feature graph. KD-DAGFM Tian et al. (2023) uses knowledge distillation and proposes a lightweight student model, directed acyclic graph FM, to learn arbitrary explicit high-order feature interactions from teacher networks. Other graph-based work, like GFM Xi et al. (2020), utilizes the popular Factorization Machine to effectively aggregate multi-order interactions in GNN. And GCFM Zheng et al. (2021) uses the multifilter graph-convolved feature crossing (GCFC) layer to learn the neighbor feature interactions.
@@
Nevertheless, GNN was originally designed for graph classification tasks, which is based on the assumption that neighbors share similar features. As a result, GNN inherits unnecessary and unsuitable operations for feature interaction modeling. Our proposed model GraphFM introduces the interaction function of FM into the neighborhood aggregation strategy of GNN to effectively capture the beneficial factorized interaction.

### v5

@@ Abstract
Factorization machine (FM) is a prevalent approach to modelling pairwise (second-order) feature interactions when dealing with high-dimensional sparse data. However, on the one hand, FMs fail to capture higher-order feature interactions suffering from combinatorial expansion. On the other hand, taking into account interactions between every pair of features may introduce noise and degrade the prediction accuracy. To solve these problems, we propose a novel approach, the graph factorization machine (GraphFM), which naturally represents features in the graph structure. In particular, we design a mechanism to select beneficial feature interactions and formulate them as edges between features. Then the proposed model, which integrates the interaction function of the FM into the feature aggregation strategy of the graph neural network (GNN), can model arbitrary-order feature interactions on graph-structured features by stacking layers. Experimental results on several real-world datasets demonstrate the rationality and effectiveness of our proposed approach. The code and data are available at https://github.com/CRIPAC-DIG/GraphCTR.
@@ 1 Introduction
Predictive analytics is a fundamental task in machine learning (ML) and data mining (DM), which involves using input features to predict an output target, such as a real value for regression or categorical labels for classification. This is particularly important for web applications, such as online advertising and recommender systems He et al. (2017); He and Chua (2017); Zhang et al. (2022). Distinct from continuous features which can be naturally found in images and audio, the features for web applications are mostly sparse and categorical. To accurately perform predictive analytics on these types of features, it is important to consider the interactions between them. As an example, consider a scenario in which we want to predict users’ preferences for movies based on five categorical variables: (1) Language = {English, Chinese, Japanese, … }, (2) Genre = {action, fiction, … }, (3) Director = {Ang Lee, Christopher Nolan, … }, (4) Stars = {Bruce Lee, Leonardo DiCaprio, … } and (5) Release Date = {1991, 1992, … }. To capture the impact of these feature interactions, a model might consider a 3rd-order cross feature such as (Genre = fiction, Director = Christopher Nolan, Starring = Leonardo DiCaprio) or (Language = Chinese, Genre = action, Starring = Bruce Lee) as potentially indicating higher user preferences.
@@
FM Rendle (2010, 2012) is a popular and effective method for modelling feature interactions, that involves learning a latent vector for each one-hot encoded feature and modelling the pairwise (second-order) interactions between them through the inner product of their respective vectors. FM has been widely used in the field of recommender systems and click-through rate predictions due to its simplicity and effectiveness. However, because FM considers all feature interactions, it has two main drawbacks.
@@
One of the main limitations of FM is that it is not able to capture higher-order feature interactions, which are interactions between three or more features. While higher-order FM (HOFM) has been proposed Rendle (2010, 2012) as a way to address this issue, it suffers from high complexity due to the combination expansion of higher-order interactions. This makes HOFM difficult to use in practice. To address the limitations of FM in capturing higher-order feature interactions, several variants have been proposed that utilize deep neural networks (DNNs) Zhang et al. (2016); He and Chua (2017); Cheng et al. (2016); Guo et al. (2017). Factorisation machine-supported neural networks (FNNs) Zhang et al. (2016) apply DNNs on top of pretrained factorization machines to model high-order interactions. Neural factorization machines (NFM) He and Chua (2017) design a bi-interaction layer to learn pairwise feature interactions and apply DNNs to learn higher-order interactions. Wide&Deep Cheng et al. (2016) introduces a hybrid architecture containing both shallow and deep components to jointly learn low-order and high-order feature interactions. DeepFM Guo et al. (2017) similarly combines a shallow component with a deep component to learn both types of interactions. While these DNN-based models can effectively learn high-order feature interactions, in an implicit, bitwise manner. Consequently, they may lack the ability to provide persuasive rationales for their outputs.
@@
In addition to not being able to effectively capture higher-order feature interactions, FM is also suboptimal because it considers the interactions between every pair of features, even if some of these interactions may not be beneficial for prediction Zhang et al. (2016); Su et al. (2020). These unhelpful feature interactions can introduce noise and lead to overfitting, as they do not provide useful information but make it harder to train the model. For example, in the context of predicting movie preferences, the feature interactions between Language and Release Date might not be relevant and, therefore, might not provide useful information for prediction. Ignoring these irrelevant feature interactions can improve model training. To solve this problem, the Attentional Decomposition Machine (AFM) Xiao et al. (2017) distinguishes the importance of the factorized interaction by reweighing each cross-feature using the attentional score Bahdanau et al. (2014), i.e. the influence of useless cross-features is reduced by assigning lower weights. However, it requires a predefined maximum order, which limits the potential of the model to find discriminative crossing features. Therefore, the Adaptive Factorization Network (AFN) Cheng et al. (2020) used a logarithmic neural transformation layer composed of multiple vector-wise logarithmic neurons to automatically learn the powers (i.e. the order) of features in a potentially useful combination, thereby adaptively learning cross-features and their weights from the data.
@@
Currently, Graph Neural Networks (GNN) Kipf and Welling (2017); Hamilton et al. (2017); Veličković et al. (2018) have recently emerged as an effective class of models for capturing high-order relationships between nodes in a graph and have achieved state-of-the-art results on a variety of tasks such as computer vision Li et al. (2017), neural language processing Marcheggiani and Titov (2017); Yao et al. (2019), and recommender systems Wang et al. (2019); Wu et al. (2019). At their core, GNNs learn node embeddings by iteratively aggregating features from neighboring nodes, layer by layer. This allows them to explicitly encode high-order relationships between nodes in the embeddings. GNNs have shown great potential for modelling high-order feature interactions for click-through rate prediction. Fi-GNN Li et al. (2019) proposed connecting each pair of features and treating the multi-field features as a fully-connected graph, using a gated graph neural network (GGNN) Li et al. (2015) to model feature interactions on the graph. Graph factorizer machine (GFM) Xi et al. (2020) utilizes FM to aggregate second-order neighbour messages, and utilizes the superposition of multiple GFM layers to aggregate higher-order neighbour messages to achieve multi-order interactions from neighborhoods for recommendation. Graph-Convolved Factorization Machines (GCFM) Zheng et al. (2021) developed the graph-convolved feature crossing (GCFC) layer to traverse all features for each input example and leveraged the features of each sample to compute the corresponding multi-feature interaction graph and propagated its influence on other features. KD-DAGFM Tian et al. (2023) proposes a directed acyclic graph based model, that can be aligned with the DP Dudzik and Veličković (2022) algorithm to improve the knowledge distillation (KD)Hinton et al. (2015) capability. However, not all feature interactions are beneficial, and GNNs rely on the assumption that neighboring nodes share similar features, which may not always hold in the context of feature interaction modelling.
@@
In summary, when dealing with feature interactions, FM suffers intrinsic drawbacks. We thus propose a novel model graph factorization machine (GraphFM), which takes advantage of the GNN to overcome the problems of FM for feature interaction modelling. By treating features as nodes and feature interactions as the edges between them, the selected beneficial feature interactions can be viewed as a graph. We thus devise a novel technique to select beneficial feature interactions, which also involves inferring the graph structure. Then, we adopt an attentional aggregation strategy to aggregate these selected beneficial interactions to update the feature representations. Specifically, to accommodate the polysemy of feature interactions in different semantic spaces, we utilize a multi-head attention mechanism Vaswani et al. (2017); Veličković et al. (2018). Each layer of our proposed model produces higher-order interactions based on the existing layers; thus, the highest-order interactions are determined by layer depth. Since our proposed approach selects the beneficial feature interactions and models them in an explicit manner, it has high efficiency in analysing high-order feature interactions and thus provides rationales for the model outcome. Through extensive experiments conducted on the CTR benchmark and recommender system datasets, we verify the rationality, effectiveness, and interpretability of our proposed approach.
@@
Overall, the main contributions of this work are threefold: (1) We analyse the shortcomings and strengths of FM and GNN in modelling feature interactions. To solve their problems and leverage strengths, we propose a novel model GraphFM for feature interaction modelling. (2) By treating features as nodes and their pairwise feature interactions as edges, we bridge the gap between the GNN and FM, and make it feasible to leverage the strength of the GNN to solve the FM problem. (3) Extensive experiments are conducted on the CTR benchmark and recommender system datasets to evaluate the effectiveness and interpretability of our proposed method. We show that GraphFM can provide persuasive rationales for feature interaction modelling and prediction-making processes.
@@ 2 Related Work
In this work, we proposed a GNN-based approach for modelling feature interactions. We design a feature interaction selection mechanism, which can be seen as learning the graph structure by viewing the feature interactions as edges between features. In this section, we review three lines of research that are relevant to this work: 1) techniques for learning feature interactions, 2) GNNs, and 3) graph structure learning methods.
2.1 Feature Interaction Modelling
@@
Modelling feature interactions is a crucial aspect of predictive analytics and has been widely studied in the literature. FM Rendle (2010) is a popular method that learns pairwise feature interactions through vector inner products. Since its introduction, several variants of FM have been proposed, including field-aware factorization machine (FFM) Juan et al. (2016) which takes into account field information and introduces field-aware embeddings, and AFM Xiao et al. (2017), which considers the weight of different second-order feature interactions. FmFM Sun et al. (2021) modelled the interactions of field pairs as a matrix and utilized a kernel product to capture field interactions. However, these approaches are limited to modelling second-order interactions, which may not be sufficient in some cases.
@@
As deep neural networks (DNNs) have proven successful in a variety of fields, researchers have begun using them to learn high-order feature interactions due to their deeper structures and nonlinear activation functions. The general approach is to concatenate the representations of different feature fields and feed them into a DNN to learn the high-order feature interactions. Factorization-machine supported Neural Networks (FNNs) Zhang et al. (2016) used pretrained factorization machines to create field embeddings before applying a DNN, while product-based neural networks (PNNs) Qu et al. (2016) model both second-order and high-order interactions through the use of a product layer between the field embedding layer and the DNN layer. Like PNNs, neural factorization machines (NFMs) He and Chua (2017) also use a separate layer to model second-order interactions, but they use a Bi-Interaction Pooling layer instead of a product layer and follow it with summation rather than concatenation. Other approaches to modelling second-order and high-order interactions jointly use hybrid architectures. The Wide&Deep Cheng et al. (2016) and DeepFM Guo et al. (2017) involve modelling low-order interactions and deep modelling high-order interactions. However, similar to other DNN-based approaches, these models learn high-order feature interactions in an implicit, bit-wise manner and may lack transparency in their feature interaction modelling process and model outputs. As a result, some studies have attempted to learn feature interactions in an explicit fashion through the use of specifically designed networks. Deep&Cross Wang et al. (2017) introduces a CrossNet that takes the outer product of features at the bit level, while xDeepFM Lian et al. (2018) uses a compressed interaction network(CIN) to take the outer product at the vector level and then compresses the resulting feature maps to update the feature representations. However, xDeepFM has been found to have issues with generalizability and scalability, and it has relatively high complexity due to its consideration of all pairwise bit-level interactions. DCNV2 Wang et al. (2021) similarly used CIN to learn efficient explicit and implicit feature intersections, but it additionally leverages low-rank techniques to approximate feature crosses in subspaces for better performance and latency trade-offs.
@@
More recently, several studies have attempted to use attention mechanisms to model feature interactions in a more interpretable way. HoAFM Tao et al. (2019) updates feature representations by attentively aggregating the representations of co-occurring features, while AutoInt Song et al. (2019) uses a multi-head self-attention mechanism to explicitly model feature interactions. InterHAt Li et al. (2020) is another model that uses an attentional aggregation strategy with residual connections to learn feature representations and model feature interactions. However, even with the use of attention mechanisms to account for the weight of each pair of feature interactions, aggregating all interactions together can still introduce noise and degrade the prediction accuracy. To address these issues, some recent studies have attempted to identify beneficial feature interactions automatically. AutoFIS Liu et al. (2020) is a two-stage algorithm that uses a gate operation to search and model beneficial feature interactions, but there is a loss of information between the stages, and the modelling process is not interpretable. AFN Cheng et al. (2020) used a logarithmic neural network to adaptively learn high-order feature interactions, and the SIGN Su et al. (2020) utilized mutual information to detect beneficial feature interactions and a linear aggregation strategy to model them. However, these approaches may not be expressive or interpretable enough.
@@ 2.2 Graph Neural Networks
A graph is a kind of data structure that reflects a set of entities (nodes) and their relationships (edges). Graph neural networks (GNNs), as deep learning architectures for graph-structured data, have attracted increasing attention. The concept of GNNs was first proposed by Gori et al. (2005), and further elaborated in Scarselli et al. (2009). Currently, most of the prevailing GNN models follow the neighbourhood aggregation strategy, that is, to learn the latent node representations by aggregating the features of neighbourhoods layer by layer. The high-order relations between nodes can be modelled explicitly by stacking layers. Gated graph neural networks (GGNN) Li et al. (2015) use GRUs Cho et al. (2014) to update node representations based on aggregated neighborhood feature information. However based on graph spectral theory Bruna et al. (2013), the learning process of graph convolutional networks (GCNs) Kipf and Welling (2017) can also be considered a mean-pooling neighborhood aggregation. GraphSAGE Hamilton et al. (2017) concatenates node features and introduces three mean/max/LSTM aggregators to pool neighborhood information. Graph attention network (GAT) Veličković et al. (2018) incorporates an attention mechanism to measure the weights of neighbors when aggregating neighborhood information of a node.
@@
Due to its strength in modelling relations on graph-structured data, the GNN has been widely applied to various applications, such as neural machine translation Beck et al. (2018), semantic segmentation Qi et al. (2017), image classification Marino et al. (2017), situation recognition Li et al. (2017), recommendation Wu et al. (2019); Chen et al. (2020); Wang et al. (2019), script event prediction Li et al. (2018), and fashion analysis Cui et al. (2019). The Fi-GNN Li et al. (2019) is the first attempt to exploit the GNN for feature interaction modelling. It first proposes connecting all the feature fields; thus, the multi-field features can be treated as a fully-connected graph. Then the GGNN Li et al. (2015) is utilized to model high-order feature interactions on the feature graph. KD-DAGFM Tian et al. (2023) uses knowledge distillation and proposes a lightweight student model, namely, directed acyclic graph FM, for learning arbitrarily explicit high-order feature interactions from teacher networks. Other graph-based work, like GFM Xi et al. (2020) utilized the popular factorization machine to effectively aggregate multi-order interactions in the GNN. In addition, GCFM Zheng et al. (2021) uses the multifilter graph-convolved feature crossing (GCFC) layer to learn the neighbor feature interactions.
@@
Nevertheless, the GNN was originally designed for graph classification tasks, and is based on the assumption that neighbors share similar features. As a result, the GNN inherits unnecessary and unsuitable operations for feature interaction modelling. Our proposed model GraphFM introduces the interaction function of FM into the neighborhood aggregation strategy of the GNN to effectively capture the beneficial factorized interaction.